OpenStudy (anonymous):

whats the quadriatic form of 16xsquared=9

OpenStudy (mathteacher1729):

\[\Huge 16x^2=9\]

OpenStudy (anonymous):

16x^2 - 9 = 0

OpenStudy (anonymous):

16x^2=25 in quadriatic form as factors

OpenStudy (anonymous):

(a-b)(a+b) = a^2 - b^2 Your equations have similar property. So, factor them.